OpenStudy (anonymous):

hi i needed help to test the convergence of the series summation( 0 to infinity) [ 2^(2n+1) / (-4)^n ]

OpenStudy (zarkon):

\[2^{2n+1}=2^{2n}2^1=4^n\cdot 2\]

OpenStudy (zarkon):

\[(-4)^n=(-1)^n4^n\]

OpenStudy (zarkon):

\[\frac{2^{2n+1}}{(-4)^n }=\frac{4^n\cdot 2}{(-1)^n4^n}=\frac{2}{(-1)^n}\]